OpenStudy (anonymous):

What is the total measure of all of the interior angles in a convex 15-gon?

OpenStudy (anonymous):

exterior angle = 360/15= 24 degrees interior angle = 180-24 = 156 degrees total measure of all angles = 15 * 156 degrees

OpenStudy (anonymous):

180(n-1) 180(15-1) 180(14) = 2520

OpenStudy (anonymous):

@jimmyrep 360 is the sum of All exterior angles OF ANY regular POLYGON. but it is not the sum of all interior angles any regular polygon. 360 is only applicable to quadrilaterals.

OpenStudy (anonymous):

i used the external angle to get interior ie 180 - ext angle . and for regular polygon total angles = n sides * int angle - thjs total is same for all n-sided polygons, regular or not.

OpenStudy (anonymous):

but you cannot use 360 for finding the interior angles of other polygons.